OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half. y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (jack1):

y - y1 = m(x - x1) is the point slope form of an equation

OpenStudy (nikato):

The third one its pretty much like the point slope formula

OpenStudy (jack1):

(x1, y1) is your known point, in this case (6, -3)

OpenStudy (anonymous):

Im never going to understand slope -.-

OpenStudy (jack1):

m is the slope, in this case 0.5 or 1/2, whichever ur more comfy with

OpenStudy (anonymous):

0.5 I dont like fractions

OpenStudy (anonymous):

so wait the answer is the 3rd one ?

OpenStudy (jack1):

ok, so use this equation: \[\large y - y_1 = m(x - x_1)\], substitute y1 for -3, x1 for 6 and m for 1/2 so whats your answer?

OpenStudy (jack1):

yep, @nikato was correct above in saying it's the 3rd one

OpenStudy (jack1):

do u understand why now tho? @Aileennx3

OpenStudy (anonymous):

No I dont understand slope period I have the hardest times with it